OpenStudy (anonymous):

Determine if the lines that pass through the given points are parallel, perpendicular or neither. Line A: (2, 3) and (--4, 5) Line B: (-4, 9) and (-3, 12)

OpenStudy (anonymous):

use the formula y2-y1/x2-x1 5-3/-4-2= 2/-6= 1/-3 12-9/ -3-(-4)= 3/1 Its perpendicular

OpenStudy (anonymous):

thanks, :)